OpenStudy (anonymous):

[48/(-6)] / (- one fourth) =

OpenStudy (anonymous):

-8/(-1/4)=(-8)(-4)=32

OpenStudy (anonymous):

LOL, thats the answer I got too but didn't think it was right. Thank you!

OpenStudy (anonymous):

\[\frac{48}{-6}\div\frac{-1}{4}\]am'I correct?